OpenStudy (anonymous):

multiply sqrt(5) * sqrt(2)=

jimthompson5910 (jim_thompson5910):

sqrt(5) * sqrt(2)= sqrt(5*2) = sqrt(10) So sqrt(5) * sqrt(2) = sqrt(10)